How much heat energy (in kj) will be absorbed or released if 11.5 grams of ammonia is produced? State whether the energy will be absorbed or released?
Answer #1

Haber process, energy released.

N2(g) + 3H2(g) <----> 2NH3(g) DH0 = -91.8 kj

2 mol NH3 formation gives 91.8 kj energy.

no of mol of NH3 = 11.5/17 = 0.676 mol

Amount of energy released = 0.676*91.8/2 = 31.03 kj